suppose you buy four buckets of apples and five buckets of peaches for $64. if you bought eight buckets of apples and three buckets of peaches, you would of spent $72. how much is one bucket of apples(a), and one bucket of peaches (p)?

Answer by jorel555(1290)

Let a=apples, p=peaches;
4a+5p=64
8a+3p=72
-----------
8a+10p=128
8a+3p=72
7p=56
p=8
Peaches cost $8/bucket, apples cost $6/bucket!!!!!!!!!!!!!!!!